# Co-Engel graphs of certain finite non-Engel groups

### v3

@@ 6. Various spectra and energies of $\mathop{\mathcal{E}_{c}^{-}}(G)$
$E(\mathop{\mathcal{E}_{c}^{-}}(G))-|v(\mathop{\mathcal{E}_{c}^{-}}(G))|$ $=2^{t+1}(m-1)-2^{t}m$
$=2^{t}(m-2)>0.$
@@
We have $|v(\mathop{\mathcal{E}_{c}^{-}}(G))|=q(p-1)$ and
$E(K_{q(p-1)})=2(q(p-1)-1).$
By Theorem 6.5 we get
$E(\mathop{\mathcal{E}_{c}^{-}}(G))-E(K_{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|})$ $=2(p-1)(q-1)-2(q(p-1)-1)$
$=-2(p-2)\leq 0.$
@@
$E(\mathop{\mathcal{E}_{c}^{-}}(G))-|v(\mathop{\mathcal{E}_{c}^{-}}(G))|$ $=2(p-1)(q-1)-q(p-1)$
$=(p-1)(q-2)>0.$
@@ 7. Zagreb indices of $\mathop{\mathcal{E}_{c}^{-}}(G)$
By Theorem 4.1 we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m\cdot 2^{t}}$. Therefore, by Lemma 7.1, we get
$M_{1}(\mathop{\mathcal{E}_{c}^{-}}(G))=2^{3t}m(m-1)^{2}.$
and
$M_{2}(\mathop{\mathcal{E}_{c}^{-}}(G))=2^{4t-1}m(m-1)^{3}.$
Here $|v(\mathop{\mathcal{E}_{c}^{-}}(G))|=2^{t}m$ and $|e(\mathop{\mathcal{E}_{c}^{-}}(G))|=2^{2t-1}m(m-1)$. Therefore
$\frac{M_{1}(\mathop{\mathcal{E}_{c}^{-}}(G))}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}=\frac{2^{3t}m(m-1)^{2}}{2^{t}m}=2^{2t}(m-1)^{2}.$
and
$\frac{M_{2}(\mathop{\mathcal{E}_{c}^{-}}(G))}{|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}=\frac{2^{4t-1}m(m-1)^{3}}{2^{2t-1}m(m-1}=2^{2t}(m-1)^{2}.$
Hence, the result follows. ∎
@@
Proof.
By Theorem 4.2 we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{q\cdot(p-1)}$. Therefore, by Lemma 7.1, we get
$M_{1}(\mathop{\mathcal{E}_{c}^{-}}(G))=q(q-1)^{2}(p-1)^{3}.$
and
$M_{2}(\mathop{\mathcal{E}_{c}^{-}}(G))=\frac{q(q-1)^{3}(p-1)^{4}}{2}.$
Here $|v(\mathop{\mathcal{E}_{c}^{-}}(G))|=q(p-1)$ and $|e(\mathop{\mathcal{E}_{c}^{-}}(G))|=\frac{q(p-1)^{2}(q-1)}{2}$. Therefore
$\frac{M_{1}(\mathop{\mathcal{E}_{c}^{-}}(G))}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}=\frac{q(q-1)^{2}(p-1)^{3}}{q(p-1)}=(q-1)^{2}(p-1)^{2}.$
and
$\frac{M_{2}(\mathop{\mathcal{E}_{c}^{-}}(G))}{|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}=\frac{q(q-1)^{3}(p-1)^{4}}{q(p-1)^{2}(q-1)}=(q-1)^{2}(p-1)^{2}.$
Hence, the result follows. ∎

### v4

@@ 7. Various spectra and energies of $\mathop{\mathcal{E}_{c}^{-}}(G)$
$E(\mathop{\mathcal{E}_{c}^{-}}(G))-|v(\mathop{\mathcal{E}_{c}^{-}}(G))|=2^{t+1}(m-1)-2^{t}m=2^{t}(m-2)>0.$
@@
We have $|v(\mathop{\mathcal{E}_{c}^{-}}(G))|=q(p-1)$ and $E(K_{q(p-1)})=2(q(p-1)-1).$ By Theorem 7.5 we get
$E(\mathop{\mathcal{E}_{c}^{-}}(G))-E(K_{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|})=2(p-1)(q-1)-2(q(p-1)-1)=-2(p-2)\leq 0.$
@@
$E(\mathop{\mathcal{E}_{c}^{-}}(G))-|v(\mathop{\mathcal{E}_{c}^{-}}(G))|=2(p-1)(q-1)-q(p-1)=(p-1)(q-2)>0.$
@@ 8. Zagreb indices of $\mathop{\mathcal{E}_{c}^{-}}(G)$
By Theorem 5.1 we have $\mathop{\mathcal{E}_{c}^{-}}(G)\cong K_{m\cdot 2^{t}}$. Therefore, by Lemma 8.1, we get the expressions for $M_{1}(\mathop{\mathcal{E}_{c}^{-}}(G))$ and $M_{2}(\mathop{\mathcal{E}_{c}^{-}}(G))$. Here $|v(\mathop{\mathcal{E}_{c}^{-}}(G))|=2^{t}m$ and $|e(\mathop{\mathcal{E}_{c}^{-}}(G))|=2^{2t-1}m(m-1)$. Therefore
$\frac{M_{1}(\mathop{\mathcal{E}_{c}^{-}}(G))}{|v(\mathop{\mathcal{E}_{c}^{-}}(G))|}=\frac{2^{3t}m(m-1)^{2}}{2^{t}m}=2^{2t}(m-1)^{2}=\frac{2^{4t-1}m(m-1)^{3}}{2^{2t-1}m(m-1}=\frac{M_{2}(\mathop{\mathcal{E}_{c}^{-}}(G))}{|e(\mathop{\mathcal{E}_{c}^{-}}(G))|}.$
Similarly, using Theorem 5.2 and Lemma 8.1, we get the following result.